OpenStudy (anonymous):

evaluate the function at the indicated variables f(x)= 3x-1; f(2x), 2f(x)

OpenStudy (anonymous):

F(x) = 3x-1 So, F(2x) = 3(2x) - 1 = 6x - 1 And, 2F(x) = 2(3x-1) = 6x - 2